OpenStudy (anonymous):

lim x-->infinity (xsin(1/x))

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} xsin(1/x)\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} (xsin(1/x))\] \[=\lim_{x \rightarrow \infty} sin(1/x)/(1/x)\] x goes to infinity = 1/x goes to 0 We know \[\lim_{x \rightarrow 0} sin (x)/x=1\] Therfore, the answer is 1.

OpenStudy (anonymous):

thx